OpenStudy (anonymous):

One leg of a right triangle is 14 inches longer than the smaller leg, and the hypotenuse is 16 inches longer than the smaller leg. Find the lengths of the sides of the triangles

OpenStudy (anonymous):

of the traingle*

OpenStudy (campbell_st):

 using Pythagoras' thorem \[x^2 +(x +14)^2 = (x +16)^2\] so \[2x^2 + 28x + 196 = x^2 + 32x + 256\] simplify then solve \[x^2 - 4x - 50 = 0\]